OpenStudy (thesecret20111):

How do I find the equation of an straight line given two points of another line that is at a perpendicular distance of 4. Points of known line are: (2,3) and (4,7)

OpenStudy (thesecret20111):



OpenStudy (science0229):

Ok. From the diagram, we can see that the 2 lines are parallel, right?

OpenStudy (science0229):

What is the equation of the line that passes through the 2 points?

OpenStudy (thesecret20111):

Thats what I understand from my question

OpenStudy (thesecret20111):

I'll work it out now

OpenStudy (thesecret20111):

I believe the equation is: y=2x-1

OpenStudy (science0229):

Good.

OpenStudy (science0229):

Because the slopes of 2 parallel line are same, the other line can be written as y=2x+b, right?

OpenStudy (thesecret20111):

Yep

OpenStudy (science0229):

Now, because the distance between 2 lines are always 4, if you pick a random point from either line and use point-line distance formula, you'll get 4.

OpenStudy (thesecret20111):

I'm not fully aware what the point-line distance formula is.. sorry :/

OpenStudy (science0229):

It's fine. Then we'll go traditional, first, then I'll teach you the shorcut.

OpenStudy (science0229):

First, let's choose a very very easy point on the 1st line since we have to find b. How does (0,b) sound to you?

OpenStudy (thesecret20111):

Yep sounds good

OpenStudy (science0229):

Ok. Now, let's say that a point on the second line is (a,2a-1).

OpenStudy (science0229):

And also say that a line (0,b) and (a,2a-1) is perpendicular to y=2x-1

OpenStudy (thesecret20111):

Ok

OpenStudy (science0229):

So the slope of (0,b) and (a,2a-1) is -1/2, right?

OpenStudy (science0229):

Also, the distance between the 2 points is 4, right?

OpenStudy (thesecret20111):

Sorry I'm a little confused as to where you got that second set of points from

OpenStudy (thesecret20111):

especially (a,2a-1)

OpenStudy (science0229):

Ok

OpenStudy (science0229):



OpenStudy (thesecret20111):

Ok so (0,b) is y-intercept of perpendicular line and (a,2a-1) is a random intersection point of perpendicular line to known line?

OpenStudy (science0229):

Exactly!

OpenStudy (thesecret20111):

Ok I get where you got those values now

OpenStudy (science0229):

Now, if I put the 2 sentences I said above into equations....\[\frac{ (2a-1)-b }{ a-0 }=\frac{ -1 }{ 2 }\]

OpenStudy (science0229):

and

OpenStudy (science0229):

\[\sqrt{(a-0)^2+((2a-1)-b)^2}=4\]

OpenStudy (science0229):

or

OpenStudy (science0229):

\[(a-0)^2+((2a-1)-b)^2=16\]

OpenStudy (science0229):

@thesecret20111 Do you know how to solve the system of these 2 equations?

OpenStudy (thesecret20111):

Do you mean solve them like simultaneous equations?

OpenStudy (science0229):

yeah.

OpenStudy (thesecret20111):

I'd need to use elimination right?

OpenStudy (science0229):

Yeah. Make sure to eliminate a since we're looking for b

ganeshie8 (ganeshie8):

it can be short if you know either `cross product` or `distance between point/line formula`, but working it this way is a good practice for ur coordinate geometry :) and also notice that there will be two lines that are equidistant from a given line in 2d

OpenStudy (science0229):

I'm planning to teach thesecret20111 the distance between point/line formula after he gets the basic idea.

OpenStudy (thesecret20111):

Sorry its been a while since I did simultaneous equations.. Do I multiply both equations by (a-0)?

OpenStudy (science0229):

For the "slope equation" solve for a in terms of b and substitute it in the "distance equation:

OpenStudy (science0229):

correction: "distance equation"

OpenStudy (thesecret20111):

I'm having trouble solving for a in the slope equation I can't see a way to get 'a' by itself

OpenStudy (science0229):

What did you do so far?

OpenStudy (science0229):

Can you show me?

OpenStudy (thesecret20111):

Its pretty bad its a=((-1/2)+(b/a)/2) +1

OpenStudy (science0229):

Ok. Let's start over. I'll show all step for isolating a

OpenStudy (science0229):

cross multiply, first.

OpenStudy (science0229):

\[-a=2(2a-1-b)\] \[-a=4a-2b-2\] \[5a=2b+2\] \[a=\frac{ 2 }{ 5 }b+\frac{ 2 }{ 5 }\]

OpenStudy (thesecret20111):

Ok I get it so far

OpenStudy (science0229):

Now substitute that in the "distance equation" to get b Remember that b has to be positive.

OpenStudy (thesecret20111):

Ok I'll give it a go. BTW I had forgotten about cross multiplying before :/

OpenStudy (science0229):

Alright.

OpenStudy (thesecret20111):

(2/5b+2/5)^2 +((2(2/5b+2/5) -1)-b)^2 = 16

OpenStudy (science0229):

Right. Keep going.

OpenStudy (science0229):

Tip: Don't foil it right away.

OpenStudy (science0229):

from the first term, factor out (2/5)^2 and from the second term, factor out (-1/5)^2

OpenStudy (thesecret20111):

Do you mean make the first term 2/5^2(b+1)?

OpenStudy (science0229):

\[(\frac{ 2 }{ 5 }b+\frac{ 2 }{ 5 })^2+(-\frac{ 1 }{ 5 }b-\frac{ 1 }{ 5 })^2=(\frac{ 2 }{ 5 }(b+1))^2+(-\frac{ 1 }{ 5 }(b+1)^2=16\]

OpenStudy (science0229):

Correction: \[(-\frac{ 1 }{ 5 }(b+1))^2\]